# Scaling Laws for Floating–Point Quantization Training

### v1

@@ 3.2 Setup
We trained and evaluated a suite of LLaMA (Dubey et al., 2024) architecture models on a subset of the Dolma V1.7 dataset (Soldaini et al., 2024), maintaining the same sampling proportion used for training the OLMo 7B-v1.7 model (Groeneveld et al., 2024). Our experiments involved a systematic exploration of language model pretraining across $N\in[41,85,154,679]$ million parameters and $D\in[10,20,50,100]$ billion tokens. Additionally, we conducted two extra pretraining sessions with models containing 1.2 billion parameters to validate the accuracy of our scaling law equation.
For each combination of $(N,D)$, we executed over 36 runs, systematically sweeping through each exponent and mantissa combination, varying the quantization target during training, and exploring different block sizes for quantization. In total, we performed 366 runs. Detailed hyperparameters and ablation studies can be found in Table 1 and Table LABEL:tab:experiments.
Hyper-parameters 41M 85M 154M 679M 1.2B
Layers 12 12 12 24 24
Hidden Size 512 768 1024 1536 2048
FFN Hidden Size 1536 2048 2816 4096 5632
Attention Heads 8 12 16 24 32
Attention Head size 64 64 64 64 64
@@
Table 1: Model hyper-parameters for each size.
3.3 Basic Scaling Law Form
Our research builds upon the foundational scaling laws for training data and model size, which are pivotal in understanding the efficiency and effectiveness of machine learning models. We first evaluate the classical scaling laws: the Chinchilla scaling law (Hoffmann et al., 2022) and the OpenAI scaling law (Kaplan et al., 2020), and then give the design of ours with precision-aware factors. To determine which scaling law better fits the empirical data, we conducted experiments using BF16 precision across a spectrum of model sizes ranging from 41 million to 679 million parameters. We plotted the resulting scaling law curves to visualize the fit between the predicted and actual losses.
(a) Chinchilla basic scaling law.
(b) OpenAI basic scaling law.
Figure 2: The fitting performance of classical scaling laws. The size of the data point is proportional to $D$.
3.3.1 Classical Scaling Law Fit
Figure 2 (left) illustrates the curve based on the Chinchilla scaling law. The empirical training losses align almost perfectly with the predicted losses, indicating a highly accurate fit. Figure 2 (right) depicts the curve based on the OpenAI scaling law. Here, the training losses also show a good match with the predicted losses, but the degree of fit is not as precise as that observed with the Chinchilla scaling law.
3.3.2 Basic Form of Our Scaling Law
Given the superior fit of the Chinchilla scaling law in our BF16 precision experiments, we proceeded to use this law as the basis for further exploration into our proposed scaling law for float precision. Our aim is to extend the understanding of scaling laws to encompass the impact of numerical precision on model performance, particularly focusing on the trade-offs between precision, computational efficiency, and model accuracy. In the following sections, we will present our methodology for investigating the scaling law for float precision, the results of our experiments, and the implications of our findings for the design and training of machine learning models at various levels of numerical precision. The marginal effects of the exponent (E), the mantissa (M), and the size of the scaling factors (B) on the performance of LLM will be discussed with experiments.
3.4 Quantization Targets
In our pursuit of balancing practicality with academic rigor, we have chosen to focus on the quantization of inputs to the General Matrix Multiplication (GEMM) computations within the Transformer architecture. Transformer consists of three main GEMM operations: forward computation, input gradient computation, and parameter gradient computation. The inputs to these matrix multiplications in both forward and backward passes include six distinct elements: $\mathbf{X}$, $\mathbf{W}$, ${\mathrm{d}\mathbf{Y}_{1}}$, ${\mathbf{W}_{\text{bwd}}}$, $\mathrm{d}\mathbf{Y}_{2}$, and $\mathbf{X}_{\text{bwd}}$, which can be quantized to ($P_{1}$) through ($P_{6}$) respectively (see Figure 3).
Figure 3: Quantization Targets. We select P2, P4, and P6 as our quantization targets for the following exploration of scaling laws.
3.4.1 Experimental Findings
Figure 3 illustrates the classical quantization targets of P1 to P6 to be explored. Through our experiments, we observed that the quantization of these inputs has varying impacts on LLM results. As illustrated in Figure 4, our key observations related to quantization targets are as follows:
  1. 1.
    The quantization of P1, P3, and P5 has a significant effect on the model’s performance, leading to a substantial increase in loss. Notably, the quantization of P5 results in a pronounced degradation of performance, with losses increasing by up to 2%. This suggests that compressing and losing information in the input embedding during the backward pass can lead to considerable performance penalties.
  2. 2.
    Quantizing only one target of P4 or P6 yields the optimal performance.
  3. 3.
    Quantizing both P2 and P6 together results in similar overall performance to quantizing P2 alone. Quantizing P2, P4, and P6 together also leads to overall results comparable to quantizing P2 alone.
Figure 4: Results of loss gaps with different quantization targets.
Figure 5: The correlations between $\gamma$,$\iota$ in Eq. (12) and $N$,$D$. $\gamma$,$\iota$ could be viewed as functions of $N$,$D$. Data point size is proportional to $D$.
3.4.2 Optimized Quantization Target
To strike a balance between efficiency and effectiveness, we have opted to quantize P2, P4, and P6. Subsequent research into float precision quantization will be based on this configuration. This decision is informed by the minimal impact on performance observed when quantizing these particular inputs, as opposed to the significant degradation caused by quantizing others. By focusing on P2, P4, and P6, we aim to maintain the integrity of the model’s performance while reaping the computational benefits of quantization.
3.5 Exponent and Mantissa
The exponent and mantissa are key components of floating-point representations. Appropriately assigning bit widths to exponent and mantissa will larger alleviate information loss in floating-point quantization. In this subsection, we attempt to discover the hidden rules between exponent/mantissa and the LLM performance.
3.5.1 Exponent
Firstly, we investigate the scaling law when Exponent serves as an independent variable. Experiments of different Exponents with various other parameter settings have been conducted, followed by our attempt of parameter fitting. It is assumed that the Exponent-related scaling law conforms more to a power-law relationship form as:
$L(E)=\frac{\gamma}{(E+0.5)^{\delta}}+\iota.$ (12)
Finally, we re-fit the data using Eq 14, obtaining the results shown in Figure 6.
Figure 6: The fitting results of our Exponent-related scaling law. Data point size is proportional to $D$.
3.5.2 Mantissa
Similar to Exponent, we also assume the Mantissa-related scaling law conforms a power-law form as:
$L(M)=\frac{\gamma^{\prime}}{(M+0.5)^{\nu}}+\iota^{\prime}.$ (15)
Jointly considering the effects of D and N in representing $\gamma^{\prime}$ and $\iota^{\prime}$, we have:
$L(N,D,M)=\frac{D^{\phi^{\prime}}}{N^{\eta^{\prime}}}\frac{1}{\gamma^{\prime}(M+0.5)^{\nu}}+\iota^{\prime}L_{BF16}.$ (16)
Surprisingly, we also find that $\phi^{\prime}\approx\beta$, $\eta^{\prime}\approx\alpha$, and $\iota^{\prime}\approx 1$. Ultimately, we adopt the form of Eq. (17) to fit the joint scaling law of Mantissa with $N$ and $D$.
The fitting result is shown in Figure 7.
Figure 7: The fitting results of our Mantissa-related scaling law. Data point size is proportional to $D$.
3.5.3 Joint Exponent & Mantissa
Integrating the scaling law results of Exponent and Mantissa when each serves as an independent variable, we can naturally organize their joint scaling law with N, D, E, M into the form of Eq. (18). The final fitting effect is presented in Figure 8.
Figure 8: The fitting results of the joint Exponent & Mantissa scaling law: Data point sizes in left, middle, and right sub-figures are proportional to $D$, $M$, and $E$, respectively.
3.6 Block Size of Scaling Factor
3.6.1 Block-wise Strategy
In the quantization process, we define the statistical range of the scaling factor as the block size (B). Since the scaling factor employs high-precision caching, when $B=1$, it is equivalent to retaining a high-precision copy of the tensor to be quantized. At this point, the model’s expressiveness should be approximately the same as that of the high-precision model: $L(N,D,B=1)\approx L_{BF16}(N,D)$. After comparing the power, linear, and logarithmic law forms, we ultimately select the following logarithmic form as the scaling law when block size serves as an independent variable:
$L(B)=\kappa\log_{2}{B}+\psi.$ (19)
We demonstrate its fitting effect on experimental data in Figure 10.
Figure 10: Our scaling law precisely forecasts validation loss for diverse block sizes. Data point sizes are directly proportional to $D$ and $B$ in the respective left and right sub-figures.
3.6.2 Channel-wise Strategy
To investigate the scaling law under the channel-wise strategy, we first utilize Eq. (20) to inversely derive the equivalent block size for different N and D cases that achieve the same validation loss as when employing the channel-wise strategy. It is found that this equivalent block size of the channel-wise strategy is approximately a constant:
$\log_{2}B_{channel}\approx 13.1567,$ (21)
which is natural since the batch size of gradient is (mostly) much larger than the hidden size ($d_{in}$ in Section 3.1.1). After incorporating this equivalent block size into Eq. (20), the fitted scenario of the channel-wise strategy is in Figure 11.
Figure 11: The fitting results of the channel-wise scaling law. The size of the data point is proportional to $D$.
3.6.3 Tensor-wise Strategy
When adopting the tensor-wise strategy, we similarly employ Eq. (20) to predict its the equivalent block size. As illustrated in Figure 12, the equivalent block size satisfies a power-law like relationship with $\frac{N}{D}$ as follows:
$\log_{2}B_{tensor}\approx\frac{N^{\omega}}{\xi D^{\eta}}.$ (22)
Likewise, substituting the power-law predicted equivalent block size into Eq. (20) yields the fitted results in Figure 13.
Figure 13: The fitting results of the tensor-wise scaling law. The size of the data point is proportional to $D$.
4 A Unified Scaling Law for Floating–Point Quantization Training
In Section 3, we have decided the basic scaling law form and quantization targets in this work, followed by detailed explorations on the marginal effects of critical factors (i.e., the exponent, the mantissa, and the block size) on LLM capabilities. In this section, we will provide the unified scaling law for floating-point quantization training, with its fitting and predictive performance as well as the insightful findings deriving from our scaling law.
4.1 The Unified Scaling Law Formation
The unified scaling law for floating-point quantization training should be able to jointly consider all factors of the data size (D), model size (N), exponent (E), mantissa (M), and block size of scaling factors (B) for precise low-precision LLM performance prediction. Based on the conclusions drawn in Eq. (2), Eq. (18), and Eq. (20), we could intuitively design the unified scaling law as follows:
where the exponent E, mantissa M, and block size B jointly represent the possible information loss of float quantized training, and $\frac{D^{\beta}}{N^{\alpha}}$ reflects, in a sense, the knowledge intensity of an LLM of size N trained on D data. Note that we could smoothly fuse the factors of E, M, and B with a unified set of hyper-parameters of $\alpha$, $\beta$, and $\gamma$.
We adopt all above 358 experiments in Section 3 containing various N, D, E, M, B settings to obtain the specific hyper-parameters in Eq. (23) and Eq. (24). To ensure the simplicity and universality of our scaling law, we pick up the Occam’s Razor to fuse or expurgate unnecessary hyper-parameters. Ultimately, the final scaling law for floating-point quantization training is articulated as follows:
where the corresponding hyper-parameters are give in Table 2. The fitting performances of Eq. (25) are given in Figure 14, which show superior capability compared to previous scaling laws in low-precision training.

### v3

@@ 3.2 Setup
We trained and evaluated a range of LLaMA (Dubey et al., 2024) architecture models on a subset of the Dolma V1.7 dataset (Soldaini et al., 2024), using the same sampling proportion as for the OLMo 7B-v1.7 model (Groeneveld et al., 2024). Our experiments systematically explored language model pretraining across $N\in\{41,85,154,679\}$ million parameters and $D\in\{10,20,50,100\}$ billion tokens. Furthermore, we conducted two additional pretraining sessions with 1.2 billion-parameter models to validate our Capybara scaling law equation. For every $(N,D)$ combination, we ran over 36 experiments, systematically varying exponents and mantissas, adjusting the quantization target during training, and exploring different block sizes for quantization. In total, we carried out 366 runs. Detailed hyperparameters and ablation studies are provided in Table 1 and Table LABEL:tab:experiments.
3.3 Basic Scaling Law Form
Our research, building on foundational scaling laws for training data and model size that are crucial for understanding machine learning models’ efficiency and effectiveness, first evaluates the classical Chinchilla (Hoffmann et al., 2022) and OpenAI (Kaplan et al., 2020) scaling laws and then presents our precision - aware design. To find which scaling law better fits empirical data, we ran experiments with BF16 precision for model sizes from 41 million to 679 million parameters and plotted scaling law curves to show the fit between predicted and actual losses. Since the Chinchilla scaling law had a better fit in our BF16 precision experiments (as shown in Figure 8), we used it as the basis for exploring our proposed float precision scaling law. Our aim is to expand the understanding of scaling laws to include the impact of numerical precision on model performance, especially focusing on the trade - offs among precision, computational efficiency, and model accuracy.
In the following, we will present our methodology for investigating float - precision scaling laws, experimental results, and implications for model design and training at different numerical precision levels. Experiments will be used to discuss the marginal effects of exponent (E), mantissa (M), and scaling factor size (B) on LLM performance.
3.4 Quantization Targets
In our pursuit of balancing practicality with academic rigor, we have chosen to focus on the quantization of inputs to the General Matrix Multiplication (GEMM) computations within the Transformer architecture. Transformer consists of three main GEMM operations: forward computation, input gradient computation, and parameter gradient computation. The inputs to these matrix multiplications in both forward and backward passes include six distinct elements: $\mathbf{X}$, $\mathbf{W}$, ${\mathrm{d}\mathbf{Y}_{1}}$, ${\mathbf{W}_{\text{bwd}}}$, $\mathrm{d}\mathbf{Y}_{2}$, and $\mathbf{X}_{\text{bwd}}$, which can be quantized to ($P_{1}$) through ($P_{6}$) respectively (see Figure 2(a)).
(a) Quantization Targets.
(b) Loss Gaps.
Figure 2: For the subsequent exploration of scaling laws, we choose P2, P4, and P6 as our quantization targets.
Experimental Findings. Figure 2(a) illustrates the classical quantization targets of P1 to P6 to be explored. Through our experiments, we observed that the FP3 quantization of these inputs has varying impacts on LLM results. As illustrated in Figure 2(b), our key observations related to quantization targets are as follows: (1) The quantization of P1, P3, and P5 has significant effects on the model’s performance, leading to a substantial increase in loss. Notably, the quantization of P5 results in a pronounced degradation of performance, with losses increasing by up to 2%. This suggests that compressing and losing information in the input embedding during the backward pass could lead to considerable performance penalties. (2) Quantizing only one target of P4 or P6 yields the optimal performance. (3) Quantizing both P2 and P6 together results in similar overall performance to quantizing P2 alone. Quantizing P2, P4, and P6 together also leads to overall results comparable to quantizing P2 alone.
Optimized Quantization Target. To balance efficiency and effectiveness, we’ve selected P2, P4, and P6 for quantization. Future research will build on this setup, chosen for its minimal impact on performance compared to other options. This approach preserves model integrity while gaining computational benefits.
3.5 Exponent and Mantissa
Exponents and mantissas are key elements in FP representations. Proper bit-width assignments for them can significantly mitigate information loss during FP quantization. Here, we aim to explore the hidden associations between exponents/mantissas and LLM performance.
Exponent. Firstly, we investigate the scaling law when Exponent serves as an independent variable. Experiments of different Exponents with various other parameter settings have been conducted, followed by our attempt of parameter fitting. It is assumed that the Exponent-related scaling law conforms more to a power-law relationship form as:
$L(E)=\frac{\gamma}{(E+0.5)^{\delta}}+\iota.$ (9)
Finally, we re-fit the data using Eq 11, obtaining the results shown in Figure 9(a).
Mantissa. Similar to Exponent, we assume the Mantissa-related scaling law conforms a power-law form as $L(M)=\frac{\gamma^{\prime}}{(M+0.5)^{\nu}}+\iota^{\prime}$. Surprisingly, jointly considering the effects of D and N in representing $\gamma^{\prime}$ and $\iota^{\prime}$, we also find that $\phi^{\prime}\approx\beta$, $\eta^{\prime}\approx\alpha$, and $\iota^{\prime}\approx 1$. Ultimately, we adopt the form of Eq. (12) to fit the joint scaling law of Mantissa with $N$ and $D$.
The fitting result is shown in Figure 9(b).
Joint Exponent & Mantissa. Integrating the scaling law results of Exponent and Mantissa when each serves as an independent variable, we can naturally organize their joint scaling law with N, D, E, M into the form of Eq. (13). The final fitting effect is presented in Figure 1(b).
3.6 Block Size of Scaling Factor
Block-wise Strategy. In the quantization process, we define the statistical range of the scaling factor as the block size (B). Since the scaling factor employs high-precision caching, when $B=1$, it is equivalent to retaining a high-precision copy of the tensor to be quantized. At this point, the model’s expressiveness should be approximately the same as that of the high-precision model: $L(N,D,B=1)\approx L_{BF16}(N,D)$. After comparing the power, linear, and logarithmic law forms, we ultimately select the following logarithmic form as the scaling law when block size serves as an independent variable:
$L(B)=\kappa\log_{2}{B}+\psi.$ (14)
The fitting effect on experimental data is shown in Figure 3.
Figure 3: Our scaling law of Eq. (15) precisely forecasts validation loss for diverse block sizes. Data point sizes are directly proportional to $D$ and $B$ in the respective left and right sub-figures.
Channel-wise Strategy. To investigate the scaling law under the channel-wise strategy, we first utilize Eq. (15) to inversely derive the equivalent block size for different N and D cases that achieve the same validation loss as when employing the channel-wise strategy. It is found that this equivalent block size of the channel-wise strategy is approximately a constant: $\log_{2}B_{channel}\approx 13.1567$, which is natural since the batch size of gradient is (mostly) much larger than the hidden size ($d_{in}$ in Section 3.1). After incorporating this equivalent block size into Eq. (15), the fitted scenario of the channel-wise strategy is in Figure 13(a).
Tensor-wise Strategy. Using the tensor-wise strategy, we apply Eq. (15) to predict the equivalent block size. As shown in Figure 12, this size follows a power-law relationship: $\log_{2}B_{tensor}\approx\frac{N^{\omega}}{\xi D^{\eta}}$. Substituting this prediction into Eq. (15) produces the fitted results in Figure 13(b).
4 A Unified Scaling Law for Floating–Point Quantization Training
In this section, we provide the unified scaling law for FP quantization training, with its fitting and predictive performance as well as the insightful findings deriving from our Capybara scaling law.
4.1 The Unified Scaling Law Formation
The unified scaling law for FP quantization training should be able to jointly consider all factors of the data size (D), model size (N), exponent (E), mantissa (M), and block size of scaling factors (B) for precise low-precision LLM performance prediction. Based on the conclusions drawn in Eq. (2), Eq. (13) and Eq. (15), we could intuitively design the unified scaling law as follows:
where the exponent E, mantissa M, and block size B jointly represent the possible information loss of FP quantized training, and $\frac{D^{\beta}}{N^{\alpha}}$ reflects, in a sense, the knowledge intensity of an LLM of size N trained on D data. Note that we could smoothly fuse the factors of E, M, and B with a unified set of hyper-parameters of $\alpha$, $\beta$, and $\gamma$.
We adopted all above 358 experiments in Section 3 containing various N, D, E, M, B settings to obtain the specific hyper-parameters in Eq. (16) and Eq. (17).
To ensure the simplicity and universality of our Capybara scaling law, we selected Occam’s Razor to fuse or expurgate unnecessary hyper-parameters. Ultimately, the final scaling law for FP quantization training is articulated as follows:
where the corresponding hyper-parameters are given in Table 2. The fitting performances of Eq. (26) are given in Figure 4, which show superior capability compared to previous scaling laws in low-precision training.
Figure 4: The fitting results of our Capybara scaling law for FP quantization training. Data point size is proportional to $D$. The star points (1.2B, 7B and 70B models) are our validation.
Furthermore, we evaluate our Capybara scaling law to predict the losses of 1.2B, 7B and 70B LLMs with different low-precision settings and trained tokens (which are viewed as our validation models that are not used in calculating hyper-parameters of our Capybara scaling law). The consistently accurate fitting results demonstrate that our Capybara scaling law performs well in larger model sizes.
4.2 Implication-1: Optimal Float Layout Analysis
Figure 5: The optimal float layouts of different bit widths.
$M_{\text{opt}}=\frac{\nu P}{\delta+\nu}-0.5.$ (19)
More details are in Appendix C. The corresponding loss scaling equation incorporates this optimal layout as:
$L(N,D,P,B)=\frac{n}{N^{\alpha}}+\frac{d}{D^{\beta}}+\epsilon+\frac{D^{\beta}}{N^{\alpha}}\cdot\frac{\log_{2}B}{\gamma_{\rho}P^{\delta+\nu}},$ (20)
where $\gamma_{\rho}=\frac{\gamma\delta^{\delta}\nu^{\nu}}{(\delta+\nu)^{\delta+\nu}}$. Figure 5 visualizes the predictive performance for different $P$, demonstrating the effectiveness of the derived layout in preserving performance under varying precisions. The optimal float layouts of FP4, FP8, and FP16 are E2M1, E4M3, and E8M7 (BF16), respectively.
4.3 Implication-2: Critical Data Size for Optimal Performance
Figure 6: Variation of loss with data size under different FP quantization settings.
From Eq. (26) and Figure 6, we can observe that there are two factors that contain $D$ and they have opposite impacts on LLM loss. Intuitively, it implies that there is an optimal data size under a certain FP quantization setting. The determination of critical data size ($D_{crit}$) stands as a critical juncture within the quantized training regimen. Upon exceeding the threshold of $D_{crit}$ with the training dataset, any additional data introduction negatively impacts the model efficacy, manifesting in a rise in validation loss instead of a decline. A comprehensive derivation for the estimation of $D_{crit}$ is delineated in Appendix D:
4.4 Implication-3: Compute-Optimality with Fixed Configurations
Fixed Data Size $D$. For a fixed $D$, the optimal precision $P_{\text{crit}}$ minimizes the loss while accounting for computational constraints. From Appendix E.1, $P_{\text{crit}}$ is expressed as:
$P_{\text{opt}}(D)=\left(\gamma_{D}D^{\beta}\log_{2}B\right)^{\frac{1}{\delta+\nu}},$ (22)
where $\gamma_{D}=\frac{\delta+\nu-\alpha}{n\alpha\gamma_{\rho}}$, which consolidates the relationships between model precision and compute efficiency. Eq. (22) suggests that we can adopt such a quantization strategy: in the early stage of training, employ aggressive quantization strategies such as FP8-E4M3, or even FP4-E2M1 which may be available in the future hardware, so as to quickly converge the model to a better level. Subsequently, as the data volume as well as the “knowledge intensity” further increase, gradually enhance the training precision to BF16, or even FP32, in order to maintain the optimal cost-effectiveness of training.
Fixed Model Size $N$. For a fixed $N$, the optimal precision depends on balancing the compute resources and maintaining the required training data size. The corresponding scaling law and derivations are in Appendix E.2. We have:
$P_{\text{opt}}(N)$ $=\left[\gamma_{N}\left(\frac{C}{k}\right)^{2\beta}N^{-(\alpha+2\beta)}\log_{2}B\right]^{\frac{1}{\delta+\nu+2\beta}}.$ (23)
$P_{\text{opt}}^{\delta+\nu+2\beta}N^{\alpha+2\beta}=\text{Constant}.$ (24)
Minimization over $N$, $D$, $P$ with Fixed Compute. Through a joint analysis of the impacts of $N$, $D$, and $P$ on the final validation loss, the relationship between cost-effective precision and expected compute budget is as:
$P_{\text{opt}}^{(\delta+\nu)\frac{\alpha+\beta}{\beta}+\alpha}=\lambda\left(\gamma_{D}\log_{2}B\right)^{\frac{\alpha+\beta}{\beta}}\left(\frac{C}{k}\right)^{\alpha},$ (25)
where $\lambda=\frac{d\beta}{n\alpha}\cdot\frac{\delta+\nu-\alpha}{\delta+\nu+\beta}$. In Appendix E.3, we present more detailed derivation processes. Based on the parameters fitted from our experimental data presented in Table 2, with the block size ($B$) set to 128 and $k=6/16$, as illustrated in Figure 7, when the total compute budget is in the range of ($10^{21}$, $10^{31}$) FP operations, the optimal cost-performance ratio precision is found to lie between 4 and 8 bits. This finding implies that training larger models with lower precision and utilizing less data yields a more cost-effective approach. Developers could rely on our implications from our Capybara scaling law to decide their optimal float-pointing quantization settings.
5 Related Work
Scaling Law of LLMs. Due to the extremely large cost of resource and time for LLM training, discovering appropriate scaling laws to accurately predict LLM capabilities under different parameters is essential for product-level training. Kaplan et al. (2020) gave the classic form of the scaling law and concludes that the performance penalty depends predictably on the ratio $N^{0.74}/D$. Hoffmann et al. (2022) modeled the final loss as a parametric function of the count of model parameters and the number of trained tokens, i.e. $E+A/N^{\alpha}+B/D^{\beta}$. Bahri et al. (2024) and Lin et al. (2024) theoretical analysis of how loss scales with the size of the training dataset and the number of parameters in a power-law manner. Previous work (Dettmers & Zettlemoyer, 2023) explored the scaling laws of LLMs with different bit precisions. Recently, Kumar et al. (2024) focused on the impact of model parameters and data volume with low precision, highlighting the possible negative impacts of more trained tokens in low-precision LLM training and serving. Other recent work (Ouyang et al., 2024) also investigated the correlations of integer quantization and LLM performance. Nevertheless, it is still not that particularly clear to conclude the scaling law for FP quantization training with respect to the specific selection of the exponent, mantissa, and block size of scaling factors.
Quantization of LLMs. The quantization technique of large language models (Lang et al., 2024; Shen et al., 2024) has received widespread attention. Xiao et al. (2023) reduced the accuracy loss during quantization by smoothing the distribution of activations and weights. Dettmers et al. (2024) combined Quantization-Aware Training and LoRA methods to implement an efficient fine-tuning method. Egiazarian et al. (2024) explored techniques for compressing large language models at very low bit rates, and Behdin et al. (2023) proposed a framework that allows each layer to be quantized independently. Although previous work (Zhang et al., 2023; Yoshida, 2023) studied the impact of exponent, mantissa and blocksize on the quantification of LLMs, the comprehensive impact of these indicators has not been systematically studied and summarized.
6 Conclusion, Limitation, and Future Work
In this work, we propose our Capybara scaling law, which functions satisfactorily as a precise guidance of future low-precision LLM training. The key factors of the data size (D), model size (N), exponent (E), mantissa (M), and block size of scaling factors (B) are carefully considered in our Capybara scaling law throughout several hundred of experiments with various precision and model settings. Besides the scaling law, we also discover some insightful implications that could instruct and enhance future FP quantization training in LLMs. We hope our findings could shed light on better low-prediction LLM training to facilitate the LLM community.
In the future, we will verify our scaling laws for FP quantization training under larger model sizes and data sizes. Currently, our explorations are conducted based on the classical Transformer architecture. Whether our scaling laws could also be smoothly applied to LLMs of other architectures (e.g., Mamba series (Dao & Gu, 2024)) is worth confirming. Moreover, our experiments focus on the classical FP quantization strategies, while other new-proposed low-bit LLM quantization methods and hardware should also be covered in the future.
Acknowledgement
Ruobing Xie is supported by the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Hyperparameter Details
The detailed hyper-parameters of our LLMs are given as follows:
Table 1: Model hyper-parameters for each size.
Hyper-parameters 41M 85M 154M 679M 1.2B 7B 70B
Layers 12 12 12 24 24 35 88
Hidden Size 512 768 1024 1536 2048 4096 8192
FFN Hidden Size 1536 2048 2816 4096 5632 10752 22016
Attention Heads 8 12 16 24 32 32 64
Attention Head size 64 64 64 64 64 64 64
@@ Appendix A Hyperparameter Details
Appendix B Fitting Details of Our Scaling Laws
We name our scaling law for Floating-point quantization training as the Capybara scaling law. Under constrained resources and space, increasing the number of capybaras can significantly reduce their survival rate and quantity once a certain density threshold is surpassed. We observe a similar phenomenon in our scaling law: with a fixed model size, expanding the data size does not consistently yield improvements when the “knowledge density” becomes too high under the pressure of low-precision training. More fitting details of our Capybara scaling laws as introduced as follows.
B.1 Numberical Fits
Our Capybara scaling law is formalized as:
$\begin{split}L(N,D,E,M,B)=\frac{n}{N^{\alpha}}+\frac{d}{D^{\beta}}+\epsilon+\frac{D^{\beta}}{N^{\alpha}}\frac{\log_{2}B}{\gamma(E+0.5)^{\delta}(M+0.5)^{\nu}},\end{split}$ (26)
where the detailed fitted constants and values are shown as follows:
Table 2: Fitted constants and their values in Eq. (26).